1) Annalisa is going to start working on mode profiling and beat note search for the old MOPA NPRO.
2) In the meantime, Manasa is working on the end table items. This will be reviewed by KA in the afternoon.
The laser at ATF is moved to the 40m when the status of 1) and 2) is determined by KA to be reasonable.
We also make the beat note measurement for the ATF laser too.
Today I installed mirrors to steer the pick-off from the main laser beam in a more free part of the PSL table and make the beat note measurement between it and the NPRO.
At the beginning I took the beam from the harmonic separator after the doubling crystal, and I was going to bring it in a less full part of the table . At the end I realized that there was already a beam steered up to a more free part of the table, and the beam is taken from the transmission of the PMC.
Tomorrow I'm going to use that beam to find the beat note with the NPRO.
I also removed almost all the steering optics that I used on the ITMY table to send the auxiliary beam for ABSL through the window parallel to the POY beam. The most important thing is that I removed the BS, which was on the same path of the POY beam (see elog 8257).